# LMU-ASC 21/25

### v1

@@ 1 Introduction
This provides a motivation for the present paper, where we construct the vertices of the space-time action for Yang-Mills theory as an integral form of world-line $n$-point functions on the supermoduli space, $\mathcal{M}$ of the $\mathcal{N}=2$ spinning particle, thus mimicking the construction of super string field theory (e.g. Witten:2012bh ; Sen:2015uaa ; Ohmori:2017wtx ). To this end, we first embed the perturbative BRST spectrum of Yang-Mills theory into the algebra of vertex operators of the world line as a state-operator map. As mentioned above, this map is not unique. We determine it by demanding that the embedding defines a quasi-isomorphism to a suitable sub-complex of $\mathcal{N}=2$ vertex operator algebra. For the quadratic space-time action derived from the world-line path integral, this amounts to “integrating in” auxiliary fields whose elimination implements the quasi-isomorphism. [^4] This is just the opposition of the common usage of homotopy transfer, where one usually integrates out fields to obtain an effective action.
@@
In the next step, we derive the cubic interaction vertex as a pull-back of the (off-shell) 3-point function to $\mathcal{M}$. This requires a choice of a Poincaré dual, $Y$ on $\mathcal{M}$ (see, e.g., Cremonini:2023ccv where this object is introduced as a cochain map between different complexes on a supermanifold), since the latter is of odd dimension two rather than one, which is the expected dimension for Yang-Mills theory Witten:2012bh ; Cremonini:2025eds . We will show that different choices of $Y$ correspond to field redefinitions in the space-time action. Furthermore, we show that there is a choice of $Y$ for which the 3-point function reproduces either the perturbative cubic interaction term, or the linear deformation of the BRST operator $Q$ in Dai:2008bh . Which object is derived depends on how picture changing is implemented on the 3rd vertex (see Witten:2012bg and Cremonini:2023ccv for a description of the picture changing operation). This is in agreement with the realization found in Cremonini:2025eds for the $\mathcal{N}=1$ world line. In addition to the standard cubic interaction, the pull-back of the path integral features additional interactions with higher form potentials that are inherently present in the vertex operator algebra, in agreement with the $\mathcal{N}=1$ world line Cremonini:2025eds . However, in $\mathcal{N}=2$ we have the possibility to project the external states onto the Fock space of the perturbative Yang-Mills multiplet. In doing so, we recover the deformed BRST operator of Dai:2008bh in a background. This establishes the equivalence of the deformation problem of $Q$ and the polynomial action derived from the world-line, which, however, is not realized as a homotopy associative algebra on a Fock space $V$ but rather a cyclic complex as in Moeller:2011zz .
@@ 2 $\mathcal{N}=2$ Spinning Particle
We will consider the BRST quantization of the $\mathcal{N}=2$ spinning particle in the Hamiltonian formulation. We recast the two worldline spinors into complex linear combinations of $\psi^{\mu}=\frac{1}{\sqrt{2}}(\psi_{1}^{\mu}+i\psi_{2}^{\mu})$, $\bar{\psi}^{\mu}=\frac{1}{\sqrt{2}}(\psi_{1}^{\mu}-i\psi_{2}^{\mu})$. After gauge fixing the Einbein $e=1$ and the world line gravitino $\chi=0$, the BRST-invariant world-line action is then given by
@@
and the world-line BRST operator is given by
@@
This model then describes the BRST-spectrum of Yang-Mills theory as a state $|\Phi\rangle$ in a suitable world-line Fock module generated by $\psi,\beta,\gamma$ and $c$ with coefficients in smooth (wave) functions. That is
@@
where the picturepicture (0,1)(0,-1) stands for the super ghost ground state $\mathrm{1}_{\psi}\delta(\bar{\gamma})$. Here, the conjugate momenta act from the left as derivatives, i.e.
@@
with $\Pi_{\mu}=p_{\mu}+A_{\mu}$ is nilpotent provided $\kappa=1$ and $A$ satisfies the non-abelian equations of motion for Yang-Mills theory. We end this review with two comments. 1) Notice that the last term in the above equation maps to 0 in $V^{(0,-1)}\mathrm{1}_{\psi}$ by the operator-state map. It can thus not be reproduced by a multi-linear map on $V^{(0,-1)}\mathrm{1}_{\psi}$. Therefore, the non-linear Yang-Mills equation cannot be written as a MC-equation on $V^{(0,-1)}\mathrm{1}_{\psi}$. 2) Here, choosing $\kappa=1$ is motivated by the desire to recover the Yang-Mills equation of motion. However, in a SFT description, this should follow from a world-line calculation. We will address these points in the following sections.
@@ 3 State-Operator Map
The first step in our construction is to identify a suitable subcomplex of the full $\mathcal{N}=2$ world-line vertex operator algebra $\left(V^{(0,-1)},Q\right)$ which contains the interaction term in (9), but encodes the same physical states as the representation (5) being quasi-isomorphic to it. We thus seek a cochain map $\iota$
@@
such that $\iota\left(V^{(0,-1)}\right)$ is quasi-isomorphic to $V^{(0,-1)}1_{\psi}$.
@@
Summing up, the previous discussion suggests that we can add to the vertex terms of the following form
@@ 4 Quadratic Action
Let us now work out the quadratic action from the world-line path integral by inserting two vertices in ghost number 0 on the world line. We have two kinds of vertices, in ghost number 0, one in the $\delta(\gamma)$ (which we will denote $\mathcal{V}^{(-1,0)}$ to indicate one $\delta$ in the $\gamma$ direction and no $\delta$ in the $\bar{\gamma}$ direction) and the other in the $\delta(\bar{\gamma})$ (which we will denote $\mathcal{V}^{(0,-1)}$) representations. The former has $R$-charge 2-2, the latter has $R$-charge +2+2. In the paragraphs above we have constructed the (first pieces of the) vertex $\mathcal{V}^{(0,-1)}$ describing the same degrees of freedom of the $\mathcal{N}=1$ theory; in order not to introduce new degrees of freedom, the vertex $\mathcal{V}^{(-1,0)}$ is constructed with the same field content, that is
@@
and the angular bracket denotes the expectation value in the world-line quantum field theory evaluated with the action given in (1). See also Cremonini:2025eds for more details. A lengthy, but otherwise straightforward calculation gives
@@ 5 Interactions
Let us now turn to the cubic interaction vertex in the space-time action, which should come from the pull-back of the world-line 3-point function to the supermoduli space $\mathcal{M}_{3}$. The introduction of the extra puncture on the world-line then naturally requires a (0|2)(0|-2)-integral form (that is, with ghost number zero and total picture number -2) on $\mathcal{M}_{3}$ which is tantamount to a double picture changing, thus producing a 2-derivative interaction [^6] Indeed, integration of a odd direction in $\mathcal{M}$ comes with an adjoint action of the supercharge on the vertex operator, see e.g. Cremonini:2025eds for more details.. This is expected for a theory of gravity, familiar from the $\mathcal{N}=(1,1)$ world sheet of the closed string, but not for Yang-Mills theory. We therefore introduce a $(0|1)$-dimensional subspace that is identified via its Poincaré dual $Y$, thus “trivialising” one of the two odd moduli we integrate over. [^7] A related construction was proposed in Berkovits:1993xq ; Berkovits:1994vy by embedding the $\mathcal{N}=1$ string in a twisted $\mathcal{N}=2$ theory. If we denote by $i:\mathcal{M}^{(0|1)}\hookrightarrow\mathcal{M}^{(0|2)}$ the embedding of the $(0|1)$-dimensional subspace in the moduli space $\mathcal{M}^{(0|2)}$, we can formally write
@@
where $i^{*}\omega$ is an integral form on the subspace. If we denote the coordinates corresponding to the two odd moduli by $\eta$ and $\bar{\eta}$, and we want to trivialise, for example, the modulus given by the $\eta$ coordinate, we can choose the representative $Y=\eta\delta(d\eta)$. Thus, the cubic term is of the form
@@
The vertex operators inserted in $\pm\infty$ have the structure (picture number, $R$-charge, etc.) as those described in the previous section. Notice that the map $Y$ has picture number 1-1, which implies that it must be paired with the rest of the integrand with total picture (on the supermoduli space) 1-1 as well. We further require $\mathcal{V}(0)$ to have zero $R$-charge if we want to interpret $S^{(3)}$ as a deformation of $Q$. The operator inserted at 0 is thus constrained to have picture number 1-1, ghost number one and $R$-charge zero. In particular, it has a different structure (in terms of $R$-charge) with respect to the vertices appearing in the kinetic term. The adjoint action by $F_{0}$ implements the pullback to the supermoduli space and will be described in the next section. Before doing so, let us, however, first comment on some properties of the Poincaré dual $Y$.
@@ 5.1 Dependence on the Poincaré Dual
The Poincaré dual $Y$ corresponds to a cohomology class identifying a certain subspace of the moduli space, as explained in the previous paragraph. Its expression depends on the choice of a representative: for example, the two choices $Y_{1}=\eta\delta(d\eta)$ and $Y_{2}=\bar{\eta}\delta(d\bar{\eta})$ are shown to correspond to the same class, as they both differ by exact terms to the same object. In particular, we have
@@ 5.2 The Middle Vertex
The counting of degrees of freedom is most easily done on-shell, where $\Box\equiv 0$. Then, (37b) implies that $\delta Z^{[2]}=0$ which by (37a) then implies Maxwell’s equation for $A^{[1]}$. It might seem natural to again impose the conventional constraint $Z^{[2]}=0$, as we did in the last section at $R$-charge 1. However, in view of the gauge invariance (38b,38e) we may also, set $A^{[3]}=0$ (effectively transferring the components from $A^{[3]}$ to $E^{[3]}$) so that $Z^{[2]}=-i\mathrm{d}{}A^{[1]}$.
@@ 5.3 Cubic Vertex
We will now obtain the cubic interaction vertex of the space-time action by integrating the pull-back over the $\mathcal{N}=2$ moduli space. The pull-back is obtained by picture changing on the $\mathcal{N}=2$ moduli space, given by the obvious extension from $\mathcal{N}=1$ Cremonini:2025eds as the finite adjoint action of
@@
That still leaves us with the choice of an integration cycle, encoded in the Poincaré dual $Y$. As there is no bosonic modulus for 3 punctures (see Cremonini:2025eds ), we only need to insert one $c$-ghost to absorb the translation zero-mode (this indeed gives us a ghost number one operator inserted in 0). Let us fix the zero $R$-charge operator as $Y=\eta\delta(d\eta)+\bar{\eta}\delta(d\bar{\eta})$.

### v2

@@ 1 Introduction
This provides a motivation for the present paper, where we construct the vertices of the space-time action for Yang-Mills theory as an integral form of world line $n$-point functions on the supermoduli space, $\mathcal{M}$ of the $\mathcal{N}=2$ spinning particle, thus mimicking the construction of super string field theory (e.g. Witten:2012bh ; Sen:2015uaa ; Ohmori:2017wtx ). However, there is an important difference: in open string theory, the moduli space of the cubic vertex is of odd dimension one, while for the $\mathcal{N}=2$ spinning particle it is of odd dimension two. Yet we need the matter content of the $\mathcal{N}=2$ world line to build a Fock space representation $V$ with a (non-injective) operator state map. For the $\mathcal{N}=1$ spinning particle described in Cremonini:2025eds there is no Fock representation $V$ on which the constraint analysis of Dai:2008bh described above can be applied. We then first embed the perturbative BRST spectrum of Yang-Mills theory into the algebra of vertex operators of the world line as a state-operator map. As mentioned above, this map is not unique. We determine it by demanding that the embedding defines a quasi-isomorphism to a suitable sub-complex of $\mathcal{N}=2$ vertex operator algebra. For the quadratic space-time action derived from the world line path integral, this amounts to “integrating in” auxiliary fields whose elimination implements the quasi-isomorphism. [^4] This is just the opposition of the common usage of homotopy transfer, where one usually integrates out fields to obtain an effective action.
@@
In the next step, we derive the cubic interaction vertex as a pull-back of the (off-shell) 3-point function to $\mathcal{M}$. This requires a choice of a Poincaré dual, $Y$ on $\mathcal{M}$ (see, e.g., Cremonini:2023ccv where this object is introduced as a cochain map between different complexes on a supermanifold), since the latter is of odd dimension two rather than one, which is the expected dimension for Yang-Mills theory Witten:2012bh ; Cremonini:2025eds . In this way, we keep the matter sector of the $\mathcal{N}=2$ world line with the corresponding Fock representation, but we are still able to describe Yang-Mills by reducing the dimension of the moduli space. We will show that different choices of $Y$ correspond to field redefinitions in the space-time action. Furthermore, we show that there is a choice of $Y$ for which the 3-point function reproduces either the perturbative cubic interaction term, or the linear deformation of the BRST operator $Q$ in Dai:2008bh . Which object is derived depends on how picture changing is implemented on the 3rd vertex (see Witten:2012bg and Cremonini:2023ccv for a description of the picture changing operation). This is in agreement with the realization found in Cremonini:2025eds for the $\mathcal{N}=1$ world line. In addition to the standard cubic interaction, the pull-back of the path integral features additional interactions with higher form potentials that are inherently present in the vertex operator algebra, in agreement with the $\mathcal{N}=1$ world line Cremonini:2025eds . However, in $\mathcal{N}=2$ we have the possibility to project the external states onto the Fock space of the perturbative Yang-Mills multiplet. In doing so, we recover the deformed BRST operator of Dai:2008bh in a background. This establishes the equivalence of the deformation problem of $Q$ and the polynomial action derived from the world line, which, however, is not realized as a homotopy associative algebra on a Fock space $V$ but rather a cyclic complex as in Moeller:2011zz .
@@ 2 $\mathcal{N}=2$ Spinning Particle
We will consider the BRST quantization of the $\mathcal{N}=2$ spinning particle in the Hamiltonian formulation. We recast the two worldline spinors into complex linear combinations of $\psi^{\mu}=\frac{1}{\sqrt{2}}(\psi_{1}^{\mu}+i\psi_{2}^{\mu})$, $\bar{\psi}^{\mu}=\frac{1}{\sqrt{2}}(\psi_{1}^{\mu}-i\psi_{2}^{\mu})$. After gauge fixing the Einbein $e=1$ and the world line gravitino $\chi=0$, the BRST-invariant world line action is then given by
@@
and the world line BRST operator is given by
@@
This model then describes the BRST-spectrum of Yang-Mills theory as a state $|\Phi\rangle$ in a suitable world line Fock module generated by $\psi,\beta,\gamma$ and $c$ with coefficients in smooth (wave) functions. That is
@@
where the picturepicture (0,1)(0,-1) stands for the super ghost ground state $\mathrm{1}_{\psi}\delta(\bar{\gamma})$. We have denoted by $\mathrm{1}_{\psi},\mathrm{1}_{c},\mathrm{1}_{\gamma}$ the ground states for which the operators $\psi,c$ and $\gamma$ act as creation operators. This means that we are considering the representation for which a state will be a polynomial in $\psi,c,\gamma$. Here, the conjugate momenta act from the left as derivatives, i.e.
@@
In particular, we have
$\bar{\psi}_{a}\mathrm{1}_{\psi}=0\ ,\ b\mathrm{1}_{c}=0\ ,\ \bar{\beta}\mathrm{1}_{\gamma}=0\ .$ (7)
@@
with $\Pi_{\mu}=p_{\mu}+A_{\mu}$ is nilpotent provided $\kappa=1$ and $A$ satisfies the non-abelian equations of motion for Yang-Mills theory. We end this review with two comments. 1) Notice that the last term in the above equation maps to 0 in $V^{(0,-1)}\mathrm{1}_{\psi}$ by the operator-state map. It can thus not be reproduced by a multi-linear map on $V^{(0,-1)}\mathrm{1}_{\psi}$. Therefore, the non-linear operatorial Yang-Mills equation cannot be reproduced as a MC-equation acting on the state $\mathrm{1}_{\psi}$. 2) Here, choosing $\kappa=1$ is motivated by the desire to recover the Yang-Mills equation of motion. However, in a SFT description, this should follow from a world line calculation. We will address these points in the following sections.
@@ 3 State-Operator Map
The first step in our construction is to identify a suitable subcomplex of the full $\mathcal{N}=2$ world line vertex operator algebra $\left(V^{(0,-1)},Q\right)$ which contains the interaction term in (10), but encodes the same physical states as the representation (5) being quasi-isomorphic to it. We thus seek a cochain map $\iota$
@@
such that $\iota\left(V^{(0,-1)}\mathrm{1}_{\psi}\right)$ is quasi-isomorphic to $V^{(0,-1)}1_{\psi}$.
@@
Here, the symbol $\frac{d}{\Box}$ represents an homotopy operator for $d^{\dagger}$. Its appearance is a consequence of the absence of $A^{[3]}$ on the left hand side of (11): by using the language of Cremonini:2025eds , we know that, in the $N=1$ scenario, we can non-locally solve the three-form $A^{[3]}$ in terms of the one-form $A^{[1]}$. This is reflected here, the moment we impose the compatibility of $Q$ with the embedding map. This non-locality is resolved by referring to the bigger sub-space of $V^{(0,1)}$ that involves new fields that vanish under the projection $\mathrm{1}_{\psi}$. In particular, the term $(dA)_{\mu\nu}$ suggests the introduction of a two-form $B_{\mu\nu}$ and the non-local term $q(\frac{d}{\Box}A)$ suggests the introduction of a three-tensor $A_{\mu\nu\rho}$ which, through the equation of motion, can be non-locally expressed in terms of $A^{[1]}$. Summing up, the previous discussion suggests that we can add to the vertex terms of the following form
@@ 4 Quadratic Action
Let us now work out the quadratic action from the world line path integral by inserting two vertices in ghost number 0 on the world line. We have two kinds of vertices, in ghost number 0, one in the $\delta(\gamma)$ (which we will denote $\mathcal{V}^{(-1,0)}$ to indicate one $\delta$ in the $\gamma$ direction and no $\delta$ in the $\bar{\gamma}$ direction) and the other in the $\delta(\bar{\gamma})$ (which we will denote $\mathcal{V}^{(0,-1)}$) representations. The former has $R$-charge 2-2, the latter has $R$-charge +2+2. In the paragraphs above we have constructed the (first pieces of the) vertex $\mathcal{V}^{(0,-1)}$ describing the same degrees of freedom of the $\mathcal{N}=1$ theory; in order not to introduce new degrees of freedom, the vertex $\mathcal{V}^{(-1,0)}$ is constructed with the same field content, that is
@@
and the angular bracket denotes the expectation value in the world line quantum field theory evaluated with the action given in (1). See also Cremonini:2025eds for more details. A lengthy, but otherwise straightforward calculation gives
@@ 5 Interactions
Let us now turn to the cubic interaction vertex in the space-time action, which should come from the pull-back of the world line 3-point function to the supermoduli space $\mathcal{M}_{3}$ of the world line with $3$ punctures. The introduction of the extra puncture on the world line then naturally requires a (0|2)(0|-2)-integral form. This is in complete analogy with string theory, where a 3-punctured sphere has no even moduli but two odd moduli corresponding to the two local SUSY transformations of the 3rd puncture that cannot be gauged away. The world line path integral then realizes the pull-back of the 3-point function to the moduli space $\mathcal{M}_{3}=\mathbb{R}^{(0|2)}$. Integration over the 2 odd moduli is then tantamount to a double picture changing, thus producing a 2-derivative interaction [^6] Indeed, as in string theory, integration of an odd direction in $\mathcal{M}$ comes with an adjoint action of the supercharge on the vertex operator, see e.g. Cremonini:2025eds for more details.. This is expected for a theory of gravity, familiar from the $\mathcal{N}=(1,1)$ world sheet of the closed string, but not for Yang-Mills theory. We therefore introduce a $(0|1)$-dimensional subspace that is identified via its Poincaré dual $Y$, thus “trivialising” one of the two odd moduli we integrate over. [^7] A related construction was proposed in Berkovits:1993xq ; Berkovits:1994vy by embedding the $\mathcal{N}=1$ string in a twisted $\mathcal{N}=2$ theory (see also Ohmori:2013zla ), where, however, the chain map property was not enforced, off shell, on the matter sector. If we denote by $i:\mathcal{M}^{(0|1)}\hookrightarrow\mathcal{M}^{(0|2)}$ the embedding of the $(0|1)$-dimensional subspace in the moduli space $\mathcal{M}^{(0|2)}$, we can formally write
@@
where $i^{*}\omega$ is an integral form on the subspace. If we denote the coordinates corresponding to the two odd moduli by $\eta$ and $\bar{\eta}$, and we want to trivialise, for example, the $\eta$-dependence of the world line correlation function $\omega$, we can choose the Poincaré dual $Y=\eta\delta(d\eta)$. While it is clear that insertion of $\eta$ eliminates the $\eta$-dependence of the superfield $\omega$, the appearance of $\delta(\mathrm{d}(\eta))$ may require some motivation: integration over $\mathrm{d}(\mathrm{d}\eta)$ is introduced to cancel the Jacobian for the change of variables form the world line gravitino $\chi$ to the the odd modulus $\eta$ after gauge fixing, i.e.
$\int[D\chi]\delta(\chi-\chi_{0})\cdots=\int\mathrm{d}\eta\;|\frac{\partial\chi}{\mathrm{d}\eta}|_{\chi_{0}}^{-1}\cdots\ .$ (33)
Then $\mathrm{d}\eta$ is the Faddeev-Popov ghost introduced to cancel this Jacobian and thus ensures independence of the choice of gauge fixing. In the present example, since we do not integrate over the $\eta$-direction of the moduli space, we set $\mathrm{d}\eta$ to zero with an extra delta function in $Y$. Thus, the cubic term is of the form
@@
The vertex operators inserted in $\pm\infty$ have the structure (picture number, $R$-charge, etc.) as those described in the previous section. Notice that the map $Y$ has picture number 1-1, which implies that it must be paired with the rest of the integrand with total picture (on the supermoduli space) 1-1 as well. We further require $\mathcal{V}(0)$ to have zero $R$-charge if we want to interpret $S^{(3)}$ as a deformation of $Q$. The operator inserted at 0 is thus constrained to have picture number 1-1, ghost number one and $R$-charge zero. Since it lives in a representation with different R-charge, it has a different structure as a multiplet with respect to the vertices appearing in the kinetic term. We will describe it in section 5.2. The adjoint action by $F_{0}$ represents the implementation of the pullback to the supermoduli space and will be described in section 5.3. Before doing so, let us, however, first comment on some properties of the Poincaré dual $Y$.
@@ 5.1 Dependence on the Poincaré Dual
The purpose of this subsection is to show that the reduction by means of a Poincaré dual is well defined and may be skipped in a first reading. The Poincaré dual $Y$ corresponds to a cohomology class identifying a certain subspace of the moduli space, as explained in the previous paragraph. Its expression depends on the choice of a representative: for example, the two choices $Y_{1}=\eta\delta(d\eta)$ and $Y_{2}=\bar{\eta}\delta(d\bar{\eta})$ are shown to correspond to the same class, as they both differ by exact terms to the same object. In particular, we have
@@ 5.2 The Middle Vertex
The counting of degrees of freedom is most easily done on-shell, where $\Box\equiv 0$. Then, (39b) implies that $\delta Z^{[2]}=0$ which by (39a) then implies Maxwell’s equation for $A^{[1]}$. It might seem natural to again impose the conventional constraint $Z^{[2]}=0$, as we did in the last section at $R$-charge 1. However, in view of the gauge invariance (40b,40e) we may also, set $A^{[3]}=0$ (effectively transferring the components from $A^{[3]}$ to $E^{[3]}$) so that $Z^{[2]}=-i\mathrm{d}{}A^{[1]}$.
@@ 5.3 Cubic Vertex
We will now obtain the cubic interaction vertex of the space-time action by integrating the pull-back over the $\mathcal{N}=2$ moduli space. The pull-back is obtained by picture changing on the $\mathcal{N}=2$ moduli space, given by the [[obvious]] extension from $\mathcal{N}=1$ Cremonini:2025eds as the finite adjoint action of
@@
As such, $\int\mathrm{d}\eta\mathrm{d}(\mathrm{d}\eta)e^{iF_{0}}=(\bar{q}+b\bar{\gamma})\delta(\bar{\beta})$ which is again familiar form string theory literature. That still leaves us with the choice of an integration cycle, encoded in the Poincaré dual $Y$. As there is no bosonic modulus for 3 punctures (see Cremonini:2025eds ), we only need to insert one $c$-ghost to absorb the translation zero-mode (this indeed gives us a ghost number one operator inserted in 0). Let us fix the zero $R$-charge operator as $Y=\eta\delta(d\eta)+\bar{\eta}\delta(d\bar{\eta})$.